OpenStudy (anonymous):

A driver was fined $20 plus $4 for every mile over the speed limit of 55 mph. If the fine was $76, how fast was the driver going?

OpenStudy (anonymous):

let x equal the driver's speed 20 + 4(x-55) = 76 20 + 4x -220 = 76 20 + 4x = 296 4x = 276 x = 69

OpenStudy (anonymous):

Thank you! :)